OpenStudy (anonymous):

complete the function table for f(x) = -2x + 1 chart x f(x) -2 ? 0 ? 1 ? 2 ?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@uri

Nnesha (nnesha):

plug x value into the equation and then solve

OpenStudy (anx):

just plug the x into the x in the equation so it looks like this for the rest f(x) = -2x+1 for 0 = -2(0) +1 = 1 for 1 = -2(1) +1 = -1 for 2 = -2(2)+1 = -3

Nnesha (nnesha):

 replace x by its value and then solve